# Measurement-based Fast Quantum State Stabilization with Deep Reinforcement Learning

### v1

@@ 5.2.3 Stabilization of the GHZ state with time delay
Figure 8: With a time delay $\mathcal{t}=0.05$ a.u., under the control of the DRL agent, 50 random initial states stably evolve to the target GHZ state. The light blue lines represent the average evolution trajectories for different initial states, and the dark blue line represents the average trajectory across all different initial states.
In order to gain a clear and informative understanding of the impact of time delay on system convergence, the following analysis was conducted:
First, the following assumptions are made:
[…]
  • A concept of “Stabilization Success Rate” is defined. The objective is to obtain convergence trajectories for $50$ different initial states, where each convergence trajectory for an initial state is constructed as an average curve of $50$ different trajectories under varying environmental noise. In other words, for $50$ different initial states, we have a total of $50*50=2500$ trajectories. To obtain these $2500$ trajectories that achieve successful convergence within the time $T_{max}$, we may need data from more than $2500$ trajectories. For instance, when there are $100$ non-convergent trajectories, the convergence success rate is calculated as $2500/(2500+100)=0.9615$.
As illustrated in Figure 9, it can be observed that with an increase in time delay, the average time required for the system state to stabilize towards the target increases, while the “Stabilization Success Rate” decreases. For instance, when the time delay is $0.05$ a.u., the average convergence time is $38.52$ a.u., with a success rate of $0.9913$. However, when the time delay is increased to $0.3$ a.u., the average convergence time extends to $91.65$ a.u., with a corresponding decrease in the success rate to $0.7909$. This aligns with common intuition, where a larger time delay leads to poorer performance of the DRL agent.
Figure 9: “Stabilization success rate” (blue) and convergence time (red) with respect to time delay for 50 random initial states stabilized to the target GHZ state under the control of the DRL agent.
6 Conclusions
In this paper, we designed a DRL agent and applied it to a quantum measurement-based feedback control problem, for stabilizing quantum entangled states. By designing an effective reward function, the trained agent can not only stabilize the quantum system to the target state with high fidelity but also exhibits good robustness. We first tested the well-trained agent in “perfect case” and compared the performance with Lyapunov-based switch method. The results showed that our approach achieves comparable fidelity in shorter time, which has the potential to reduce the interaction between quantum systems with their environment thereby significantly reducing additional noise issues arising from prolonged system-environment interactions. We then tested our agent in different “imperfect cases” including measurement inefficiency and time delay in the feedback loop, where our proposed agent showed good robustness. Lastly, due to the generality of our proposed control framework, extending it to stabilize other quantum states poses minimal challenges. We believe that exploring the behaviors that may exist in real-world environments will be an interesting and challenging research direction. In the future, we will also consider reducing the traning time of DRL agents by designing more efficient and faster algorithms.
[The Impact of Different Reward Functions on Agent Performance] We take the two problems described in Section 5 (stabilization of 2- and 3-qubit systems) and train our agents using the following three different reward functions to compare their impact on the agents’ performance.
  1. 1.
    Sparse Reward: Only rewards are given when the agent reaches specific states. For example, in games like Go, it is challenging to set rewards for every intermediate move, and the state space is vast, so rewards are only obtained when the final victory is achieved. In our quantum system environment, we can set a reward when the distance $D_{\rho_{t}}\leq 0.001$. At all other times, the reward is 00.
  2. 2.
    Linear Reward: As shown in the Figure 10, the system’s reward is linearly related to the distance $D_{\rho_{t}}$.
  3. 3.
    IPF Reward: Refer to Section 4.2 for details.
Figure 10: Linear Reward Function
To ensure fair comparisons, for a target state, we maintain consistent training parameters for all three agents, with the only difference being the reward functions used during training. Subsequently, we test these trained agents using $50$ random initial states. As depicted in Figure 11(a), in the context of the 2-qubit state stabilization problem, agents trained with linear reward (orange) and IPF reward (blue) perform almost identically, both being able to stabilize the system to the desired state within a reasonable time frame. However, the sparse reward function (green) demonstrates a clear inadequacy, as the agent only receives rewards when reaching specific target states, hindering its ability to effectively explore and learn during the training process. Moving on to the more complex 3-qubit system, as shown in Figure (b), the agent trained with the sparse reward function struggle to stabilize the system to the target state. Conversely, the agent trained with the linear reward function exhibit good performance and the ability to reach the target state. However, in comparison to the IPF reward function, the agent guided by the linear reward function lack speed in achieving stability. In contrast, our proposed IPF reward function excels in stabilizing the target state. It provides the agent with balanced guidance, enabling efficient learning and faster convergence to the desired stable states. These highlight the effectiveness and superiority of the IPF reward approach in improving agent performance.
Figure 11: The performance of agents trained with three different reward functions in (a) two-qubit system (b) three-qubit (GHZ) system.

### v2

@@ V-B3 Stabilization of the GHZ state with time delay
Figure 9: With a time delay $\mathcal{t}=0.05$ a.u., under the control of the DRL agent and the Lyapunov-based method, $50$ random initial states stably evolve to the target GHZ state. The light blue lines represent the average evolution trajectories for different initial states, and the dark blue line represents the average trajectory across all different initial states.
VI Analyzing the Advantages of PNR Reward Function
In this section, the 3-qubit example will be used to analyze the advantages of the PNR reward function. As described in Section IV-B, the PNR employs a partition parameter $d$ to divide the state space into two regions: the Proximity Zone ($D_{\rho_{t}}<d$) and the Exploration Zone ($D_{\rho_{t}}\geq d$). The reward function in each region follows the nonlinear form specified in (7).
The parameters for the PNR are set as follows:
  • $d=0.001$,  $\mathfrak{e}=2$,  $\mathfrak{f}=10$;
  • $\overline{R}_{P}=100$,  $\underline{R}_{P}=1$,  $\overline{R}_{E}=0$,  $\underline{R}_{E}=-0.1$.
To evaluate the effectiveness of the proposed approach, we consider several alternative reward functions for training DRL agent. For fair comparisons, all agents are trained with the same parameters for a given target state, with the only difference being the reward functions used during training. The purpose of the analysis is to give us a clear and comprehensive understanding to the impact of the various modules of the designed reward function on the control performance, as well as to compare its performance with existing reward function designs.
The following assumptions are made:
  • The maximum evolution time of the system, denoted as $T_{\text{max}}$, is set to $100$ a.u.. For any trajectory where the system does not converge to the target state within $T_{\text{max}}$ (i.e., $D_{\rho_{T_{\text{max}}}}>0.001$), it is considered a non-convergent trajectory. For the purpose of the follow-up calculation of the average time to convergence, the “convergence time” for these non-converging trajectories up to $100$ a.u. is set to $100$ a.u..
  • A metric termed the Stabilization Success Rate is defined. During testing, $50$ distinct initial states are considered for stabilization to the target, with each initial state being stabilized under $50$ different environmental noise conditions. This results in a total of $50\times 50=2500$ individual trajectories. The Stabilization Success Rate is calculated as the ratio of the number of convergent trajectories within $T_{\text{max}}=100$ to the total number of trajectories ($2500$).
VI-A Reward Function Designs
The reward functions tested are categorized as follows:
  • Partitioned Nonlinear Reward 1 (PNR1): This follows the structure of PNR, with the only difference being the parameters $\mathfrak{e}=10$ and $\mathfrak{f}=2$, such that the slope decreases as the distance diminishes.
  • Partitioned Linear Reward (PLR): This design mirrors the partitioned structure of PNR but employs a linear reward function in each region instead of a nonlinear one. The bounds remain consistent with PNR.
  • Partitioned Sparse Reward (PSR): [16, 17] use a sparse reward structure that provides rewards only at the last time step. We consider similar design ideas, retaining the partition structure for consistency:
    • In the Proximity Zone, the reward is a fixed positive constant.
    • In the Exploration Zone, no reward is assigned.
  • Non-Partitioned Nonlinear Reward (NPNR): This reward design does not use partitioning of the state space. Common reward function designs, such as those in [21, 22], apply a uniform reward structure across all states, treating the entire state space equally without distinguishing between different regions. The NPNR reward function adopts the nonlinear form described in (7), defined over the entire state space $D_{\rho_{t}}\in[0,1]$. The rewards are non-positive, ranging from $[-1,0]$, where the penalty magnitude decreases as $D_{\rho_{t}}$ gets closer to 00.
  • Non-Partitioned Linear Negative Reward (NPLNR): Similar to NPNR, this design does not partition the state space. The reward function is linearly defined over $D_{\rho_{t}}\in[0,1]$ and is always non-positive $[-1,0]$. The penalty decreases linearly as $D_{\rho_{t}}$ approaches 00.
  • Non-Partitioned Linear Positive Reward (NPLPR): The only distinction between this design and NPLNR is that the non-partitioned reward values are strictly positive, ranging from $[0,100]$. The reward increases linearly as $D_{\rho_{t}}$ approaches 00.
  • Fidelity-Based Positive Reward (FPR): This non-partitioned reward is based on fidelity, a common approach in machine learning for quantum measurement feedback problems[21, 22]. Specifically, we adopt the reward function proposed in [22]:
    $r=F(t)^{4}+4\cdot F(t)^{25},$ (18)
    where $F(t)$ is the fidelity at time $t$. The reward is positive and increases as $D_{\rho_{t}}$ approaches 00.
VI-B Performance Comparison
To facilitate comparison, the reward functions are categorized into three types. Our PNR serves as the baseline for evaluating and comparing these categories.
VI-B1 Category 1: Partitioned Nonlinear Rewards
This category contains PNR1. The effect of the parameters $\mathfrak{e}$ and $\mathfrak{f}$ on the performance of the agent is evaluated. Using $50$ random initial states, each averaged over $50$ trajectories, the results in Fig. 10 show that PNR1 performs worse than PNR. The average convergence time for PNR1 is $12.19$ a.u., compared to $10.41$ a.u. for PNR. Additionally, the convergence success rate for PNR1 is $99.2\%$, lower than the $100\%$ success rate of PNR. These results suggest that setting $\mathfrak{e}$ smaller than $\mathfrak{f}$ is more effective, meaning the reward function’s steepness increases as the target is approached.
Figure 10: The effect of $\mathfrak{e}$ and $\mathfrak{f}$ parameters in the PNR reward function on the performance of the DRL algorithm. PNR: $\mathfrak{e}=2$ and $\mathfrak{f}=10$; PNR1: $\mathfrak{e}=10$, $\mathfrak{f}=2$.
VI-B2 Category 2: Partitioned Linear and Sparse Rewards
This class includes PLR and PSR, whose only difference from our PNR is that the form of their reward functions are linear and sparse rewards, respectively. As shown in Fig. 11, PLR achieves relatively good performance, with a convergence time of $11.80$ a.u. and a success rate of $99.76\%$, although it still underperforms compared to PNR. This highlights the superiority of our designed nonlinear reward function. In contrast, PSR performs poorly, as sparse rewards alone are insufficient for complex systems.
Figure 11: Effect of reward function on DRL algorithm under partitioning conditions. PNR: partitioned nonlinear reward; PLR: partitioned linear reward; PSR: partitioned sparse reward.
VI-B3 Category 3: Non-Partitioned Rewards
This category evaluates NPNR, NPLNR, NPLPR, and FPR. These non-partitioned reward functions include linear, nonlinear, and fidelity-based, and the reward values are either positively encouraging (gradually increasing the positive reward value) or negatively spurring (gradually decreasing the negative reward value). Fig. 12 shows that among these designs, NPNR and NPLNR perform the best with convergence times of $13.60$ a.u. and $18.48$ a.u. and success rates of $99.88\%$ and $99.56\%$, respectively. These results suggest that decreasing negative penalties near the target is more effective than increasing positive rewards. In addition, the outperformance of NPNR over NPLNR highlights the role of nonlinear reward functions. The commonly used FPR performs poorly, with a convergence time and success rate of $43.59$ a.u. and $97.72\%$, respectively, highlighting its limitations in stabilizing complex quantum states. NPLPR’s method of incrementing the positive rewards is nearly ineffective, which coincides with Remark $1$.
Figure 12: Effect of DRL algorithms with non-partitioned reward functions when not partitioned. NPNR: nonlinear negative reward; NPLNR: linear negative reward; NPLPR: linear positive reward; FPR: fidelity-based reward.
Our PNR outperforms other reward designs, demonstrating superior performance in stabilizing complex quantum states like the GHZ state. Nonlinear reward functions, especially when combined with partitioning, are shown to be more effective than linear or sparse reward structures in achieving faster convergence and higher success rates.
Table I summarizes the parameters and characteristics of these reward function designs, as well as their performance.
TABLE I: Comparison of Reward Functions
Reward Type Partitioned $\mathfrak{e},\mathfrak{f}$ Reward Function PZ Rewards EZ Rewards NP Rewards Time (a.u.) Success Rate
PNR Yes $\mathfrak{e}=2,\mathfrak{f}=10$ Nonlinear [1, 100] [-0.1, 0] \ 10.41 100%
PNR1 Yes $\mathfrak{e}=10,\mathfrak{f}=2$ Nonlinear [1, 100] [-0.1, 0] \ 12.19 99.2%
PLR Yes \ Linear [1, 100] [-0.1, 0] \ 11.8 99.76%
PSR Yes \ Sparse 1 0 \ 53.25 62.08%
NPNR No \ Nonlinear \ \ [-1, 0] 13.6 99.88%
NPLNR No \ Linear \ \ [-1, 0] 18.48 99.56%
NPLPR No \ Linear \ \ [0, 100] 97.23 3.08%
FPR No \ Fidelity-Based \ \ Fidelity 43.59 97.72%
PZ (EZ) Rewards: Reward values in Proximity (Exploration) Zone;  NP Rewards: Reward values when not partitioned.
VII Conclusions
In this work, we designed a DRL agent and applied it to measurement-based feedback control for stabilizing quantum entangled states. With a designed reward function, the trained agent can quickly stabilize the quantum system to the target state with high fidelity and demonstrates strong robustness.
The DRL agent was tested under a “perfect case” and its performance was mainly compared to a Lyapunov-based switching method, and several other DRL algorithms with different reward functions. The results showed that our approach achieve comparable fidelity in a shorter time, potentially mitigating noise caused by prolonged interactions. The agent was also evaluated under “imperfect cases,” including low measurement efficiency and feedback delay, where it maintained great performance under these challenging conditions.
To analyze our method, we conducted ablation studies on the reward function to identify the role of each component in achieving the final performance. We also compared our reward function with fidelity-based designs from the literature and several variants of our reward function, showing the superior performance of our approach.
The proposed DRL-based framework has the potential to be extended to stabilize any multi-qubit system. Naturally, it is expected that as the system dimension increases, both the training time and the time required for state estimation using the SME will grow. Future research could explore methods to reduce the dimensionality of the state space and enhance the speed of quantum state estimation based on the SME. Furthermore, developing control strategies that rely on less system information, such as using only measurement currents, may allow for efficient stabilization of the target state with sufficient fidelity. These directions aim to further improve the practical applicability of the proposed control framework.
[PPO algorithm] In the field of DRL, PPO algorithm has become a prominent method due to its effectiveness and stability in policy optimization. This appendix is dedicated to elucidating the core algorithmic concepts of PPO and providing a detailed description of its mechanisms. The aim is to provide the reader with a comprehensive understanding of how the algorithm works and facilitate its application in practical scenarios.The core of PPO is to maximize the expected cumulative returns by optimizing the strategy parameters while ensuring that the updates remain stable. The following paragraphs delve into the specifics of these algorithmic ideas.
-A Core Algorithmic Ideas of PPO
To maximize $J(\theta)$, which indicates that our chosen policy parameters $\theta$ can lead to higher average rewards, we adopt the well-known gradient descent method. Thus, we take the derivative of the expected reward $J(\theta)$ in (20), resulting in the expression shown in
$\nabla J(\theta)$ $=\sum_{\tau}R(\tau)\nabla p_{\theta}(\tau)$
$=\sum_{\tau}R(\tau)p_{\theta}(\tau)\nabla\log\,p_{\theta}(\tau)$
We use $\nabla f(x)=f(x)\nabla\log f(x)$ to derive the second row of (21). The last approximate equation is a result of practical gradient computations, where instead of calculating the expected reward for an entire trajectory, rewards contributed by each individual state-action pair ($\rho,a$) are computed separately. These individual rewards are then summed up to obtain the total cumulative reward for the optimization process. The direction of the policy update $\pi_{\theta}(a|\rho)$ is biased towards favoring state-action pairs that contribute to higher cumulative rewards within the sequence. For instance, if an action $a$ executed in state $\rho$ leads to a positive cumulative discounted reward, the subsequent update will enhance the probability of choosing action $a$ in state $\rho$, while diminishing the likelihood of selecting other actions. The update equation for the parameters $\theta$ is as follows:
$\theta=\theta+\eta\nabla J(\theta),$ (22)
where $\eta$ is the learning rate.
Once the policy $\pi_{\theta}(a|\rho)$ is updated, it necessitates the reacquisition of training data prior to the subsequent policy update. This arises due to the alteration in the probability distribution $p_{\theta}(\tau)$ brought about by the modified policy. Following data sampling, the parameter $\theta$ undergoes refinement, leading to the discarding of all prior data. Subsequent parameter updates mandate the collection of fresh data, constituting the fundamental principle underlying the conventional PG algorithm. However, in the context of quantum systems, the process of sampling system information is often characterized by time-intensive and computationally demanding operations. For instance, after each measurement, a classical computer is requisitioned to solve the SME (1) to ascertain the system’s state in the subsequent moment. This inability to reutilize previously acquired data contributes to a protracted training process. To address this challenge, an additional strategy $\pi_{\theta^{\prime}}$ is introduced, mirroring the architecture of $\pi_{\theta}(a|\rho)$. Instead of directly engaging with the environment for data gathering, the primary agent $\pi_{\theta}$ employs the auxiliary agent $\pi_{\theta^{\prime}}$ to interact with the environment and accumulate data. The objective is to subsequently utilize this data to train $\pi_{\theta}$ multiple times, effectively reducing the computational and resource demands for data collection. Ensuring the consistency of data sampled by $\pi_{\theta^{\prime}}$ with that of $\pi_{\theta}$, importance sampling [42] is introduced to facilitate this synchronization process. This approach contributes to enhancing data reuse and the overall efficiency of the training procedure. (21) is updated as:
$\nabla J(\theta)$
Here, all the state-action pairs $(\rho_{t},a_{t})$ (or alternatively, all trajectories $\tau\in\mathbb{D}$) are sampled from $\pi_{\theta^{\prime}}$. The term ${p_{\theta}(\rho_{t},a_{t})}/{p_{\theta^{\prime}}(\rho_{t},a_{t})}$ represents the importance weight, which dynamically adjusts the weighting of data sampled by $\pi_{\theta^{\prime}}$ in real time to more accurately estimate the expected value under the target policy $\pi_{\theta}$.
The corresponding objective function from (23) can be calculated as:
$J_{\rm{PPO}}^{\theta^{\prime}}(\theta)$
$J_{\rm{PPO}}^{\theta^{\prime}}(\theta)$
  • One set of main policy parameters $\theta$, which is updated every time.
-B Generalized parameter selection for training agents
Neural Network Architecture: Each policy $\pi_{\theta}$ is represented by a neural network that maps a given state $\rho$ to a probability distribution over actions $a$. The action distribution is modeled as a gaussian distribution. The input layer is processed by two fully connected hidden layers, each with $128$ neurons, accompanied by a linear output layer with the same dimension as the action space ($2$ in this paper). All hidden layers use the Tanh activation function. The value function $V^{\phi}(\rho_{t})$ is composed of a similar neural network architecture, with the only difference being that the output layer is a single linear unit used to estimate the state-value function. The value function $V^{\phi}(\rho_{t})$ is estimated using the temporal difference (TD) method [43]. Then, the generalized advantage estimator (GAE) [44] is employed to compute the advantage function in (8), which is subsequently used in (25) to calculate the gradient for updating the policy $\pi_{\theta}$.